OpenStudy (anonymous):

What is the change in entropy for a system that has 250 J of heat added to it at 25°C? -10 J/K, -0.84 J/K, 0.84 J/K, or 10 J/K **thank you!!

OpenStudy (michele_laino):

it is simple, we have to apply this formula: \[\Large \Delta S = \frac{Q}{T}\] where \delta S is the entropy change

OpenStudy (anonymous):

okay!! so 250/25 = 10 ? so our solution is 10 J/K ?

OpenStudy (michele_laino):

no, since T is the absolute temperature

OpenStudy (anonymous):

ohhh i am unsure then :/

OpenStudy (michele_laino):

we have: \[T = t + 273\] t is the Celsius temperature

OpenStudy (anonymous):

ohh so we would get 273+25=298?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

250/298=0.84? so our solution is 0.84 J/K?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

yay! thank you!! :D

OpenStudy (michele_laino):

thank you!! :D